The five number summary of a dataset was found to be:

46, 56, 60, 67, 70

An observation is considered an outlier if it is below:

An observation is considered an outlier if it is above:
Answer #1

From the five number summery,

Min , Q1 , Q2 , Q3 , max = 46 , 56 , 60 , 67 , 70

Q1 = 56 , Q3 = 67

InterquartiIe range QR = Q3 - Q1 = 67 - 56 = 11

An observation is considered an outlier if it is below Q1 - 1.5 IQR

= 56 - 1.5 * 11

= 39.5

An observation is considered an outlier if it is above Q3 +  1.5 IQR

= 67 + 1.5 * 11

= 83.5